the time in seconds required to produce a P.D at 20v across a capacitor at 1000uF when is at a steady rate of 200uC/sec

the  time in seconds required to produce a P.D at 20v across a capacitor at 1000uF when is at a steady rate of 200uC/sec

Grade:12

1 Answers

Sher Mohammad IIT Delhi
10 years ago
q=cv
(dq/dt)t=cv
t=cv/dq/dt
=1000/200=5 seconds
sher mohammad
faculty askiitians
b.tech, iit delhi